# Free Majorana Fermion Meets Gauged Ising Conformal Field Theory on the Fuzzy Sphere

### v2

@@ 2 Model
In the setup of the fuzzy sphere [66], we consider one flavour of fermions $f$ and one flavour of bosons $b$ with the same electric charge on the sphere under the influence of a magnetic monopole at its centre. We let the monopole charge they experience differ by $2\pi$, i. e., we let the fermion experience a $4\pi q$ monopole and the bosons experience a $4\pi(q-\tfrac{1}{2})$ monopole. As will be elaborated below, the mismatch of the monopole flux reflects the coupling between the electric charge and the spatial curvature, and it also allows the fermionic local operators like $f^{\dagger}b$ to have half-integer angular momenta. Due to the presence of the monopole, the single-particle eigenstates form highly degenerate quantised Landau levels [67]. The lowest Landau level (LLL) has angular momentum $q$ and degeneracy $N_{mf}=2q+1$ for the fermions and angular momentum $q-\tfrac{1}{2}$ and degeneracy $N_{mb}=2q$ for the bosons. By setting the single particle gap the leading energy scale, we project the system onto the LLL and express the fermion and boson operators in terms of the creation and annihilation operators on the LLL
$f^{\dagger}(\mathbf{r})=\frac{1}{R}\sum_{m=-q}^{q}f^{\dagger}_{m}Y_{qm}^{(q)}(\mathbf{r}),\qquad b^{\dagger}(\mathbf{r})=\frac{1}{R}\sum_{m=-q+1/2}^{q-1/2}b^{\dagger}_{m}Y_{q-1/2,m}^{(q-1/2)}(\mathbf{r}),$ (2)
where the radius of the sphere is set as $R=N_{mf}^{1/2}$ [23], and the single particle wavefunctions are the spin-weighted spherical harmonics [68].
@@
    The local electric charge density interaction $H_{U}=\int\mathrm{d}^{2}\mathbf{r}\,(n_{f}(\mathbf{r})+n_{b}(\mathbf{r}))^{2}$, where $n_{f}(\mathbf{r})=f^{\dagger}(\mathbf{r})f(\mathbf{r})$, $n_{b}(\mathbf{r})=b^{\dagger}(\mathbf{r})b(\mathbf{r})$. This term is a natural form of interaction that exists in many other fuzzy sphere models (e. g. Refs. [23, 26, 46, 49]).
@@
$H_{t}$ $=\sum_{\{m_{i}\}}M_{m_{1}m_{2}m_{3}m_{4}}^{(t)}b^{\dagger}_{m_{1}}b^{\dagger}_{m_{2}}f_{m_{3}}f_{m_{4}}+\text{h.c.}$
The four-fermion contact terms are equivalently expressed in terms of the pseudopotentials [67, 23]
$M_{m_{1}m_{2}m_{3}m_{4}}^{(t)}$ $=\frac{\delta_{m_{1}+m_{2},m_{3}+m_{4}}}{2q-1}\begin{pmatrix}q-\tfrac{1}{2}&q-\tfrac{1}{2}&2q-1\\ m_{2}&m_{1}&-m_{1}-m_{2}\end{pmatrix}\begin{pmatrix}q&q&2q-1\\ m_{3}&m_{4}&-m_{3}-m_{4}\end{pmatrix}$
$M_{m_{1}m_{2}m_{3}m_{4}}^{(b)}$ $=\frac{\delta_{m_{1}+m_{2},m_{3}+m_{4}}}{2q-1}\begin{pmatrix}q-\tfrac{1}{2}&q-\tfrac{1}{2}&2q-1\\ m_{2}&m_{1}&-m_{1}-m_{2}\end{pmatrix}\begin{pmatrix}q-\tfrac{1}{2}&q-\tfrac{1}{2}&2q-1\\ m_{3}&m_{4}&-m_{3}-m_{4}\end{pmatrix}$
$M_{m_{1}m_{2}m_{3}m_{4}}^{(bf)}$ $=\frac{\delta_{m_{1}+m_{2},m_{3}+m_{4}}}{2q-\tfrac{1}{2}}\begin{pmatrix}q&q-\tfrac{1}{2}&2q-\tfrac{1}{2}\\ m_{2}&m_{1}&-m_{1}-m_{2}\end{pmatrix}\begin{pmatrix}q&q-\tfrac{1}{2}&2q-\tfrac{1}{2}\\ m_{3}&m_{4}&-m_{3}-m_{4}\end{pmatrix}$.
@@
$H_{t}=\int\mathrm{d}^{2}\mathbf{r}\,(b^{\dagger}f)(\mathbf{r})(D_{\theta}+iD_{\phi})(b^{\dagger}f)(\mathbf{r})+\text{h.c.}$ (4)
where $D^{\mu}$ is the covariant derivative on the sphere. [^5] Specifically, $D^{\mu}=\partial^{\mu}+iqA^{\mu}$, where $b^{\dagger}f$ experiences a monopole flux q=1/2q=-1/2, and $A^{\phi}=\operatorname{ctg}\theta$ Interestingly, this pair conversion can be understood as a kinetic term of the Majorana fermion (cf. Section 5).
Let us consider some limiting cases for $\mu$. At large negative $\mu$, fermions are energetically favourable and the fermion orbitals are fully filled, forming a fermionic integer quantum Hall (fIQH) state with $\nu_{f}=1$. At large positive $\mu$, bosons are energetically favourable, and the boson density interaction drives a bosonic Pfaffian (bPf) state [54, 69]. The mismatch of number of orbitals and particles $N_{e}=N_{mb}+1$ is in agreement with the Wen-Zee shift required by the bPf state [70, 71] which arises from the coupling between the electric charge and the spatial curvature. We will discuss this phase in more detail in Secion 3.
@@ 3 Phase Diagram
We study this model numerically through exact diagonalisation (ED) using our numerical package FuzzifiED [72]. [^6] The module Fuzzifino, which is part of FuzzifiED, specialises in the ED for boson-fermion mixed systems. Sample codes can be found at FuzzifiED.jl/examples/free_majorana_spectrum.jl and free_majorana_correlator.jl. The maximal system size we reach is $N_{mf}=13$, and the maximal RAM usage is 24 gigabytes. We first study the phase diagram on the $t$-$\mu$ plane. For finite $t$, we find that the fermion density drops continuously from $1$ to 0 with increasing $\mu$ (Figure 1). [^7] For $t=0$, the number of fermions and bosons are separately conserved, level crossings appear for states in different $N_{f}$ sectors. The positive and negative $t$’s give exactly the same results, as they are connected by a $2\pi$-rotation along $z$ axis.
Figure 1: The phase diagram on the $\mu$-$t$ plane. The colour denotes the average fermion density calculated at $N_{mf}=12$. Selected contours are marked by dashed lines as visual guidance. The circles separate the fermionic integer quantum Hall (fIQH) phase and the $f$-wave chiral topological superconductor (TSC), and the squares separate TSC and bosonic Pfaffian (bPf) phase. The approximate phase boundary is determined through the scaling of the energy gap from $N_{mf}=12$ and $10$ data.
To study the nature of the fIQH-bPf transition, we first fix $t=0.3$. We calculate the finite size behaviour of the fermionic gap $\Delta E_{f}$, defined as the energy difference between the lowest excited state with angular momentum $l=1/2$ and the ground state at $l=0$, and the bosonic gap $\Delta E_{b}$ defined as the energy difference between the two lowest $l=0$ states. At a Lorentz-invariant critical point, the gap is supposed to close as $\Delta E\propto 1/R$. We find two such critical points. At $\mu_{c1}\approx 0$, both bosonic and fermionic gap close; at $\mu_{c2}\approx 2.3$, the bosonic gap closes while the fermionic gap remains (Figure 2). We keep the data only for even $N_{mf}$ near $\mu_{c2}$ because the bPf ground state is only compatible with even $N_{mf}$; for odd $N_{mf}$, the phase and the transition still exist, but the states have an anyon excitation (cf. Section 6). Similar scenario holds for other finite $t$. For each $t$, we determine the approximate phase boundary by comparing the energy gaps for $N_{mf,1}=12$ and $N_{mf,2}=10$ (Figure 1): We minimise $\Delta E(N_{mf,1})/\Delta E(N_{mf,2})](N_{mf,1}/N_{mf,2})^{1/2}$ respectively for the fermion gap and for the boson gap to get $\mu_{c1}$ and $\mu_{c2}$. The minima are typically value close to $1$.
@@
The existence of two gapless points indicates that there is an intermediate phase whose transitions to fIQH and bPf phases are both continuous. To determine the nature of this intermediate phase, we calculate the real space entanglement spectrum with the entanglement cut at the equator [73, 74]. For a 3D gapped topological phase, the low-lying part of the ground-state entanglement spectrum for a local bipartition is generally believed [75] to reproduce the spectral counting of the physical edge CFT of that phase. In the following, we analyse the entanglement spectrum of the three phases respectively.
Figure 3: The real space entanglement spectrum with the cut at the equator at different phases at different charge sector $Q=(Q_{e,A}-Q_{e,B})/2$: (a) the fermionic integer quantum Hall phase at $\mu=-2$, (b) the $f$-wave chiral topological superconductor phase at $\mu=1$ and (c) the bosonic Pfaffian phase at $\mu=5$. Markers are plotted with transparency, so that nearly degenerate levels appear darker. The degeneracies for different charge sectors are labelled beneath. In the calculation we take $N_{mf}=10$.
@@
$f$-wave chiral topological superconductor (TSC)
We continue with the analysis of the intermediate phase. We shall show that this phase is an $f$-wave TSC. The $f$-wave TSC can be regarded as a $p$-wave TSC on the background of an fIQH state. In the language of Kitaev’s ‘16-fold way’ classification of topological superconductors from spinless fermion [76], the $f$-wave TSC has the index $\nu_{K}=3$ (Chern number, or number of chiral Majorana fermions on the edge). [^8] In this language, the $p$-wave TSC has $\nu_{K}=1$, and the fIQH can be regarded as a $d$-wave TSC with $\nu_{K}=2$. An intuition for its origin is the the fluctuating number of fermions in this phase and the pairing pattern given by $H_{t}$ which antisymmetrises the orbital index of two fermions.
The TSC is described by a $\mathrm{SO}(3)_{1}$ CS theory. The 2D chiral CFT on its edge is three free chiral Majorana fermions. They form a vector under a $\mathrm{SO}(3)$ global symmetry on the edge, of which a $\mathrm{U}(1)$ subgroup is manifest on the fuzzy sphere as the charge conservation. In the entanglement spectrum, we observe a well-split lowest chiral edge mode whose degeneracy is consistent with the 2D chiral CFT (Figure 3b, Eq. (15)). The TQFT has a chiral central charge $c_{-}=3/2$. The enhancement from $\mathrm{U}(1)_{e}$ to the $\mathrm{SO}(3)$ global symmetry on the edge can be seen from the formation of multiplets. The excited states at $m=1/2$ and $1$ form $\mathrm{SO}(3)$-vectors corresponding to the Majorana fermion operator and the symmetry current.
@@
The large positive $\mu$ belongs to the bosonic Pfaffian phase. The bosonic Pfaffian state is given by the many-body wavefunction on a plane with complex coordinates $z,\bar{z}$ [54]
@@
where $\operatorname{Pf}$ is the Pfaffian of an even-dimensional square matrix, which also restricts that the bPf state is only compatible with even number of particles, i. e. $N_{mf}\in 2\mathbb{Z}$. It is described by a Chern-Simons theory $\mathrm{SU}(2)_{2}$. As $\mathrm{SU}(2)$ is a double covering of $\mathrm{SO}(3)$, $\mathrm{SO}(3)_{1}\cong\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}$, i. e., the $\mathrm{SU}(2)_{2}$ can be regarded as gauging the $\mathbb{Z}_{2}$ fermion parity symmetry from $\mathrm{SO}(3)_{1}$. Correspondingly, its chiral edge mode is the same as removing the $m\in\mathbb{Z}+\tfrac{1}{2}$ levels and keeping the $m\in\mathbb{Z}$ part of the chiral edge mode in $\mathrm{SO}(3)_{1}$ (Figure 3c).
@@
Here $\mathbb{I}$ is the vaccuum, $\psi$ is a fermionic anyon and $\sigma$ is a non-Abelian anyon. The anyons are also referred to as 0, $\tfrac{1}{2}$ and $1$. This TO can also be understood as a Ising topological order stacked with an Abelian topological order
@@
The stacking with $\mathrm{U}(1)_{4}$ adds $1$ to the chiral central charge and add an additional phase $\pi/4$ to the $\sigma$ anyon compared with their values $c_{-}^{(\text{Ising})}=1/2$ and $\theta^{\text{(Ising)}}_{\sigma}=\pi/8$ in the Ising TO. In Kitaev’s ‘16-fold way’ classification [76], this phase is described by the gauged TSC with index $\nu_{K}=3$.

### v3

@@ 2 Model
In the set-up of the fuzzy sphere [69], we consider one flavour of fermions $f$ and one flavour of bosons $b$ with the same electric charge on the sphere under the influence of a magnetic monopole at its centre. We let the monopole charge they experience differ by $2\pi$, i. e., we let the fermion experience a $4\pi q$ monopole and the bosons experience a $4\pi(q-\tfrac{1}{2})$ monopole. As will be elaborated below, the mismatch of the monopole flux reflects the coupling between the electric charge and the spatial curvature, and it also allows the fermionic local operators like $f^{\dagger}b$ to have half-integer angular momenta. Due to the presence of the monopole, the single-particle eigenstates form highly degenerate quantised Landau levels [70]. The lowest Landau level (LLL) has angular momentum $q$ and degeneracy $N_{mf}=2q+1$ for the fermions and angular momentum $q-\tfrac{1}{2}$ and degeneracy $N_{mb}=2q$ for the bosons. By setting the single-particle gap the leading energy scale, we project the system onto the LLL and express the fermion and boson operators in terms of the creation and annihilation operators on the LLL
$f^{\dagger}(\mathbf{r})$ $=\frac{1}{R}\sum_{m=-q}^{q}f^{\dagger}_{m}Y_{qm}^{(q)}(\mathbf{r}),$ $b^{\dagger}(\mathbf{r})$ $=\frac{1}{R}\sum_{m=-q+1/2}^{q-1/2}b^{\dagger}_{m}Y_{q-1/2,m}^{(q-1/2)}(\mathbf{r}),$ (2)
where the radius of the sphere is set as $R=N_{mf}^{1/2}$ [25], and the single-particle wave-functions are the spin-weighted spherical harmonics [71]. [^7] An alternative choice without a flux mismatch is to let the bosons carry the two component spinor spherical harmonics $\Omega^{(q)}_{q-1/2,qm}$ [72] with spin-weight $q$ instead of the scalar spherical harmonics $Y^{(q-1/2)}_{q-1/2,m}$ with spin-weight q1/2q-1/2. These two choices assert the same spatial profile $\|\Omega^{(q)}_{q-1/2,qm}(\mathbf{r})\|^{2}=|Y^{(q-1/2)}_{q-1/2,m}(\mathbf{r})|^{2}$ and differ only by a local basis transformation of the spinors. The equivalence shows that the physical intepretation of the flux mismatch is the coupling of an additional spin-$1/2$ degree of freedom.
@@
    The local electric charge density interaction $H_{U}=\int\mathrm{d}^{2}\mathbf{r}\,(n_{f}(\mathbf{r})+n_{b}(\mathbf{r}))^{2}$, where $n_{f}(\mathbf{r})=f^{\dagger}(\mathbf{r})f(\mathbf{r})$, $n_{b}(\mathbf{r})=b^{\dagger}(\mathbf{r})b(\mathbf{r})$. This term is a natural form of interaction that exists in many other fuzzy-sphere models (e. g. Refs. [25, 38, 35, 47]).
@@
$H_{t}$ $=\sum_{\{m_{i}\}}M_{m_{1}m_{2}m_{3}m_{4}}^{(t)}b^{\dagger}_{m_{1}}b^{\dagger}_{m_{2}}f_{m_{3}}f_{m_{4}}+\text{h.\penalty 10000\ c.}$
The four-fermion contact terms are equivalently expressed in terms of the pseudo-potentials [70, 25]
$M_{m_{1}m_{2}m_{3}m_{4}}^{(t)}$ $=\delta_{m_{1}+m_{2},m_{3}+m_{4}}(4q-1)\begin{pmatrix}q-\tfrac{1}{2}&q-\tfrac{1}{2}&2q-1\\ m_{2}&m_{1}&-m_{1}-m_{2}\end{pmatrix}\begin{pmatrix}q&q&2q-1\\ m_{3}&m_{4}&-m_{3}-m_{4}\end{pmatrix}$
$M_{m_{1}m_{2}m_{3}m_{4}}^{(b)}$ $=\delta_{m_{1}+m_{2},m_{3}+m_{4}}(4q-1)\begin{pmatrix}q-\tfrac{1}{2}&q-\tfrac{1}{2}&2q-1\\ m_{2}&m_{1}&-m_{1}-m_{2}\end{pmatrix}\begin{pmatrix}q-\tfrac{1}{2}&q-\tfrac{1}{2}&2q-1\\ m_{3}&m_{4}&-m_{3}-m_{4}\end{pmatrix}$
$M_{m_{1}m_{2}m_{3}m_{4}}^{(bf)}$ $=\delta_{m_{1}+m_{2},m_{3}+m_{4}}\,4q\begin{pmatrix}q&q-\tfrac{1}{2}&2q-\tfrac{1}{2}\\ m_{2}&m_{1}&-m_{1}-m_{2}\end{pmatrix}\begin{pmatrix}q&q-\tfrac{1}{2}&2q-\tfrac{1}{2}\\ m_{3}&m_{4}&-m_{3}-m_{4}\end{pmatrix}$.
@@
$H_{t}$ $=\int\mathrm{d}^{2}\mathbf{r}\,\eta(\mathbf{r})D_{+}\eta(\mathbf{r})+\text{h.\penalty 10000\ c.},$ $\eta(\mathbf{r})$ $=b^{\dagger}(\mathbf{r})f(\mathbf{r}),$ $D_{\pm}=D_{\theta}\pm iD_{\phi}$ (4)
where $D_{\mu}$ is the covariant derivative on the sphere. [^8] Specifically, $D^{\mu}=\partial^{\mu}+iqA^{\mu}$ and $A^{\phi}=\operatorname{ctg}\theta$. The derivatives act as the raising and lowering operators the spin-weighted spherical harmonics $D_{\pm}Y_{lm}^{(s)}=\pm\sqrt{(l\mp s)(l\pm s+1)}Y_{lm}^{(s\pm 1)}$ The operator $\eta=b^{\dagger}f$ experiences a monopole flux q=1/2q=-1/2. This pair conversion can be understood as a kinetic term of the Majorana fermion (cf. Section 7).
Let us consider some limiting cases for $\mu$. At large negative $\mu$, fermions are energetically favourable and the fermion orbitals are fully filled, forming a fermionic integer quantum Hall (fIQH) state with $\nu=1$. At large positive $\mu$, bosons are energetically favourable, and the boson density interaction drives a bosonic Pfaffian (bPf) state [58, 73]. The mismatch of number of orbitals and particles $N_{e}=N_{mb}+1$ is in agreement with the Wen-Zee shift required by the bPf state [74, 75] which arises from the coupling between the electric charge and the spatial curvature. We will discuss this phase in more detail in Secion 3.
@@ 3 Phase Diagram
We study this model numerically through exact diagonalisation (ED) using our numerical package FuzzifiED [76]. [^9] The module Fuzzifino, which is part of FuzzifiED, specialises in the ED for boson-fermion mixed systems. Sample codes can be found at FuzzifiED.jl/examples/free_majorana_spectrum.jl, free_majorana_correlator.jl, and super_ising_spectrum.jl. The maximal system size we reach is $N_{mf}=13$, and the maximal RAM usage is 24 gigabytes. We first study the phase diagram on the $t$-$\mu$ plane. For finite $t$, we find that the fermion density drops continuously from $1$ to 0 with increasing $\mu$ (Figure 1). [^10] For $t=0$, the number of fermions and bosons are separately conserved, level crossings appear for states in different $N_{f}$ sectors. The positive and negative $t$’s give exactly the same results, as they are connected by a $2\pi$-rotation along $z$ axis.
Figure 1: The phase diagram on the $\mu$-$t$ plane. The colour denotes the average fermion density calculated at $N_{mf}=12$. Selected contours are marked by dashed lines as visual guidance. The circles separate the $\nu=1$ fermionic integer quantum Hall (fIQH) phase and the $\nu_{K}=3$ Majorana quantum Hall (MQH) phase, and the squares separate MQH and bosonic Pfaffian (bPf) phase. The approximate phase boundary is determined through the scaling of the energy gap from $N_{mf}=12$ and $10$ data.
To study the nature of the fIQH-bPf transition, we first fix $t=0.3$. We calculate the finite-size behaviour of the fermionic gap $\Delta E_{f}$, defined as the energy difference between the lowest excited state with angular momentum $l=1/2$ and the ground state at $l=0$, and the bosonic gap $\Delta E_{b}$ defined as the energy difference between the two lowest $l=0$ states. At a Lorentz-invariant critical point, the gap is expected to close as $\Delta E\propto 1/R$. We find two such critical points. At $\mu_{c1}\approx 0$, both bosonic and fermionic gap close; at $\mu_{c2}\approx 2.3$, the bosonic gap closes while the fermionic gap remains (Figure 2). We keep the data only for even $N_{mf}$ near $\mu_{c2}$ because the bPf ground state is only compatible with even $N_{mf}$; for odd $N_{mf}$, the phase and the transition still exist, but the states have an anyon excitation (cf. Section 6). Similar scenario holds for other finite $t$. For each $t$, we determine the approximate phase boundary by comparing the energy gaps for $N_{mf,1}=12$ and $N_{mf,2}=10$ (Figure 1): We minimise $\Delta E(N_{mf,1})/\Delta E(N_{mf,2})](N_{mf,1}/N_{mf,2})^{1/2}$ respectively for the fermion gap and for the boson gap to get $\mu_{c1}$ and $\mu_{c2}$. The minima are typically value close to $1$.
@@
The existence of two gapless points indicates that there is an intermediate phase whose transitions to fIQH and bPf phases are both continuous. To determine the nature of this intermediate phase, we calculate the real-space entanglement spectrum with the entanglement cut at the equator [77, 78]. For a 3D gapped topological phase, the low-lying part of the ground-state entanglement spectrum for a local bi-partition is generally believed [79] to reproduce the spectral counting of the physical edge CFT of that phase. In the following, we analyse the entanglement spectrum of the three phases respectively.
Figure 3: The real-space entanglement spectrum with the cut at the equator at different phases at different charge sector $Q=(Q_{e,A}-Q_{e,B})/2$: (a) the fermionic integer quantum Hall phase at $\mu=-2$, (b) the $\nu_{K}=3$ Majorana quantum Hall phase at $\mu=1$ and (c) the bosonic Pfaffian phase at $\mu=5$. Markers are plotted with transparency, so that nearly degenerate levels appear darker. The degeneracies for different charge sectors are labelled beneath. In the calculation we take $N_{mf}=10$.
@@
Majorana quantum Hall phase (MQH)
We continue with the analysis of the intermediate phase. We shall show that this phase is a MQH phase with Kitaev index $\nu_{K}=3$ [57] (the $\nu=1$ fIQH can be viewed as $\nu_{K}=2$ MQH). The $\nu_{K}=3$ MQH is described by a $\mathrm{SO}(3)_{1}$ CS theory. The 2D chiral CFT on its edge is three free chiral Majorana fermions. They form a vector under a $\mathrm{SO}(3)$ global symmetry on the edge, of which a $\mathrm{U}(1)$ subgroup is manifest on the fuzzy sphere as the charge conservation. In the entanglement spectrum, we observe a well-split lowest chiral edge mode whose degeneracy is consistent with the 2D chiral CFT (Figure 3b, Eq. (26)). The TQFT has a chiral central charge $c_{-}=3/2$. The enhancement from $\mathrm{U}(1)_{e}$ to the $\mathrm{SO}(3)$ global symmetry on the edge can be seen from the formation of multiplets. The excited states at $m=1/2$ and $1$ form $\mathrm{SO}(3)$-vectors corresponding to the Majorana fermion operator and the symmetry current. An intuition for its origin is the the fluctuating number of fermions in this phase and the pairing pattern given by $H_{t}$ which anti-symmetrises the orbital index of two fermions.
@@
The large positive $\mu$ belongs to the bosonic Pfaffian phase. The bosonic Pfaffian state is given by the many-body wave-function on a plane with complex co-ordinates $z,\bar{z}$ [58]
@@
where $\operatorname{Pf}$ is the Pfaffian of an even-dimensional square matrix, which also restricts that the bPf state is only compatible with even number of particles, i. e. $N_{mf}\in 2\mathbb{Z}$. It is described by a Chern-Simons theory $\mathrm{SU}(2)_{2}$. As $\mathrm{SU}(2)$ is a double covering of $\mathrm{SO}(3)$, $\mathrm{SO}(3)_{1}\cong\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}$, i. e., the $\mathrm{SU}(2)_{2}$ can be regarded as gauging the $\mathbb{Z}_{2}$ fermion-parity symmetry from $\mathrm{SO}(3)_{1}$. Correspondingly, its chiral edge mode is the same as removing the $m\in\mathbb{Z}+\tfrac{1}{2}$ levels and keeping the $m\in\mathbb{Z}$ part of the chiral edge mode in $\mathrm{SO}(3)_{1}$ (Figure 3c).
@@
Here $\mathbb{I}$ is the vacuum, $\psi$ is a fermionic anyon and $\sigma$ is a non-Abelian anyon. The anyons are also referred to as 0, $\tfrac{1}{2}$ and $1$. This TO can also be understood as an Ising topological order stacked with an Abelian topological order
@@
The stacking with $\mathrm{U}(1)_{4}$ adds $1$ to the chiral central charge and add an additional phase $\pi/4$ to the $\sigma$ anyon compared with their values $c_{-}^{(\text{Ising})}=1/2$ and $\theta^{\text{(Ising)}}_{\sigma}=\pi/8$ in the Ising TO. In Kitaev’s ‘16-fold way’ classification [57], this phase is described by the gauged MQH with index $\nu_{K}=3$.